OpenStudy (anonymous):

log base 2 of (x+5) + log base 2 of (x+4)= log base 2 of 2

OpenStudy (anonymous):

since the base of log 2 is uniform i just omit writing it hence the question goes log(x+5) +log(x+4)=log2 thus log(x+5)(x+4)=log2 taking antilog both sides wrt common base 2 as mentioned we have (x+5)(x+4)=2 x^2+9x+20-2=0 x^2+9x+18=0 x^2+6x+3x+18=0 x(x+6)+3(x+6)=0 or (x+6)(x+3)=0 hence x=-3,-6 but putting x=-6 in the question we have log(x+5)=log(-1) (absurd) hence x=-3

OpenStudy (anonymous):

rem log a+ log b=logab (provided base of log is uniform throughout)

OpenStudy (anonymous):

but if you plug -6 and -3 back in you get log 2